When Monique rented a boat to a cruise on the Sandy River, she was told that the boat takes 1.5 times as long to go 360 miles up a river than to return. If the boat cruises at 15 miles per hour in still water, what is the rate of the current?
Answer by uma(370)

Let the rate of current be x mph
Speed of the boat = 15 mph
Distance covered = 360 miles
Speed while going up the river = 15 - x mph
Speed while going down the river = 15 + x mph
Time taken for upward journey = 360/(15 - x)
Time taken for downward journey = 360/(15 + x)
Thus the equation is 360/(15-x) = (1.5) [360/(15+x)]
==> 1/(15-x) = 1.5/(15+x) [dividing by 360 on both sides]
==> 15+x = 1.5(15-x)
==> 15+x = 22.5 - 1.5x
==> x + 1.5x = 22.5 - 15
==> 2.5x = 7.5
==> x = 7.5/2.5
==> x = 3

Thus the rate of the current = 3 mph
Good Luck!!!